How long does it take (in minutes) for light to reach the Earth from the Sun, a distance of 1.586 x 10^8?
1
Expert's answer
2018-03-19T08:42:22-0400
Solution: the speed of light is c=3×108 m/s . The distance to be overcome, according to the condition of the problem, is l=1.586×108 km. Then the time will be: t=l/c=(1.586×〖10〗^11)/(3×〖10〗^8 )=530 s. In one minute 60 seconds, then 530 seconds will be: 530/60=8 min⁡〖42 second〗
Answer: 8 min 42 second